# Crossing the Functional Desert: Critical Cascades and a Feasibility Transition for the Emergence of Life

### v1

@@ Abstract
The origin of life poses a problem of combinatorial feasibility: How can persistent functional organization arise in exponentially branching assembly spaces when unguided exploration behaves as a memoryless random walk? We show that nonlinear threshold-cascade dynamics in connected interaction networks provide a minimal, substrate-agnostic mechanism that can soften this obstruction. Below a critical connectivity threshold, cascades die out locally and structured input–output response mappings remain sparse and transient—a “functional desert” in which accumulation is dynamically unsupported. Near the critical percolation threshold, system-spanning cascades emerge, enabling persistent and discriminative functional responses. We illustrate this transition using a minimal toy model and generalize the argument to arbitrary networked systems. Also near criticality, cascades introduce structural and functional persistence, directional bias, and weak dynamical path-dependence into otherwise memoryless exploration, allowing biased accumulation over short coherence timescales. This connectivity-driven transition—functional percolation—requires only generic ingredients: interacting units, nonlinear thresholds, influence transmission, and non-zero coherence times. The mechanism does not explain specific biochemical pathways, but it identifies a necessary dynamical regime in which structured functional organization can emerge and persist, providing a physical foundation for how combinatorial feasibility barriers can be crossed through network dynamics alone.
@@
Figure 1: Even a microscopic collection of interacting units can support functional accumulation. Minimal interaction structure already supports non-trivial functional responses. Two input nodes AA and $B$ (e.g., coarse-grained stimuli or activation sites) influence a downstream node $C$ through directed couplings. Node $C$ is a thresholded unit with threshold parameter $\phi$: It activates if the weighted fraction of active inputs exceeds $\phi$. For $\phi\leq 0.5$, activation of either input suffices and $C$ implements an OR-like response. For $\phi>0.5$, both inputs are required and $C$ implements an AND-like response. Because the interaction links and response rule persist over the cascade timescale, structured multi-input responses can recur over short coherence intervals within the same dynamical regime, supporting a minimal form of functional accumulation.
1 Introduction
A central challenge in the origin of life is the emergence of persistent functional organization in vast combinatorial spaces. Previous work has shown that unguided exploration behaves effectively as a memoryless random walk, making reliable accumulation dynamically unsupported rather than merely improbable [4, 5]. In such spaces, the number of possible molecular or functional configurations grows super-exponentially with system size, rendering reliable accumulation of organized structure vanishingly unlikely without mechanisms that bias or stabilize exploration. Here, accumulation refers not to the growth of specific molecular objects, but to a more generalized notion: The emergence and persistence of structured functional organization, as revealed through non-trivial system responses to environmental inputs. In this sense, our usage aligns with Schrödinger’s qualitative characterization of life as maintaining and amplifying order in far-from-equilibrium conditions [13].
@@ 1 Introduction
Thus, while prior work has emphasized the infeasibility of assembling complex structures in vast configuration spaces, the present framework makes clear that the deeper obstruction lies in function space: Below a critical interaction regime, structured input–output mappings cannot persist. Therefore the above ’feasibility desert’ becomes a ’functional desert’, in which meaningful functional organization is dynamically unsupported.
This motivates the central question of the present work: Can cascade dynamics provide a minimal physical mechanism that relieves the combinatorial feasibility obstruction? Specifically, we ask whether connectivity-driven transitions in cascade behavior can enable the emergence of persistent, structured functional responses in otherwise unguided systems, thereby allowing accumulation in exponentially branching spaces.
@@ 3.1 Toy Example: Minimal Functional Cascades
Imagine three system elements—molecules, neurons, cells, or other interacting units—that come together randomly and establish effective channels of influence among them. If these interactions are sufficient to form a minimally connected component, then activity at one element can, in principle, reach the others. In this informal sense, the system has crossed a basic connectivity threshold analogous to percolation: Local interactions are no longer isolated, but linked into a coherent interaction structure.
@@
Importantly, such functional responses require only a minimal form of persistence. The interaction links must remain intact long enough for influence to propagate, and the response rule at C must remain coherent long enough for inputs to be integrated. No long-term memory, storage, or repeatability is assumed. The persistence required is strictly local in time and limited to the duration of the cascade itself.
In this sense, even a tiny connected system can begin to “cross the functional desert.” Rather than requiring the accumulation of specific molecular objects, the system accumulates functional organization: Structured input–output relationships that can, in principle, be reused and extended. The toy model provides a concrete, intuitive demonstration of how minimal interactions can already support nontrivial functional structure.
3.2 General Mechanism: Cascades and Persistence
The toy example illustrates a more general mechanism that applies to networked systems of arbitrary size. Consider a system composed of many interacting nodes. As connectivity increases, the system undergoes a percolation transition: A giant connected component emerges, allowing influences to propagate across large portions of the network [17].
@@ 3.2 General Mechanism: Cascades and Persistence
Under these minimal conditions, cascades can integrate multiple inputs into organized system-level responses. Functional organization emerges directly from connectivity and nonlinear dynamics, without requiring pre-specified functions or stored programs. This connectivity-driven transition in the accessibility of functional responses, coinciding with cascade emergence near percolation criticality, is what we refer to as functional percolation, and has been studied elsewhere [17, 18, 19].
@@ 4 Consequences of the General Mechanism
4.1 Persistence
The core argument implies two minimal and distinct forms of persistence. Structural persistence refers to the coherence of interaction connectivity and susceptibility profiles over the timescale of a cascade. For influence to propagate, the network structure must remain intact long enough for signals to traverse multiple nodes [17, 6].
Functional persistence refers to the coherence of response mappings over the realization timescale. For a system to exhibit a structured functional response, the rules governing how nodes integrate inputs must remain stable long enough for inputs to be transformed into organized outputs [18, 14].
Crucially, neither form of persistence assumes long-term stability, stored memory, inheritance, or replication. Persistence here means only that relevant structures remain coherent over the duration of a single realization. This minimal notion of persistence is sufficient to support cascades and functional responses without invoking biological machinery or evolutionary processes [19].
@@ 4.2 Accumulation
In purely memoryless exploration of combinatorial spaces, partial progress is lost between realizations, preventing reliable buildup of organization, much like the ”monkeys typing Shakespeare” intuition [3]. Structural and functional persistence prevent this total erasure by allowing coherent response patterns to recur across perturbations.
@@ 4.3 Directionality and Bias
In this framework, directionality refers not to the direction of signal propagation per se, but to the directionality of accumulation in functional and assembly space [5]. Cascades, together with minimal structural and functional persistence, bias which functional organizations are realized and can recur over short coherence timescales.
Because node responses depend on collective inputs, certain response patterns recur more readily than others. This introduces a bias in how structured functional organizations are explored and stabilized, even in the absence of design, optimization, or selection [17, 12].
@@
Importantly, this directionality is non-teleological. It does not reflect purpose, intention, or evolutionary selection, but arises as a structural consequence of cascades operating in persistent interaction networks [8].
@@ 5 Feasibility and Crossing the Functional Desert
Figure 2: Functional percolation as a feasibility transition in function space (schematic). As effective interaction connectivity increases, threshold-cascade dynamics organize into three regimes. Below the transition (subcritical), cascades die out locally and realizable input–output response functions are sparse and nonpersistent. Near the functional percolation threshold, non-saturating system-spanning cascades emerge, sharply expanding the accessible repertoire of persistent response functions and enabling long-range directed influence. Above the transition (supercritical), increasing connectivity strengthens collective coupling, reducing functional diversity and constraining responses toward more correlated global patterns without eliminating information flow. The near-critical regime constitutes a narrow feasible window in which accumulation in an exponentially branching space becomes possible through persistence and bias, without collapse or saturation of functional distinctions.
@@
Near the critical transition, however, cascades are system-spanning without being fully saturating. This regime supports maximal functional discrimination, long-range directed influence, and the emergence of a rich repertoire of persistent response mappings [Fig. 2] [19, 18]. Only in this narrow window does structured functional organization become both realizable and retainable across realizations. An ’oasis’ in the functional desert.
Thus, near criticality, cascades bias which functional organizations are realized. Exploration becomes directionally constrained in functional and assembly space, introducing path-dependence in what is built [5, 14]. Accumulation is no longer symmetric or transient, but structured and persistent. Quantitative characterization of this weak dynamical path-dependence in specific models is left for future work.
In this sense, functional percolation constitutes a feasibility transition. Random exploration in combinatorial spaces is converted into directionally biased accumulation through the emergence of persistent, discriminative cascades [5, 14]. Accumulation becomes physically supported, not merely statistically possible.
@@ 6 Discussion
Across these domains, structured functional responses to inputs correspond to collective behaviors with non-zero persistence across realizations, enabled by interaction topology, response thresholds, and timescale coherence rather than by detailed biochemical specialization.
Empirical assembly-theory studies emphasize that real chemical systems occupy a highly structured and historically constrained region of object space, characterized by reuse of substructures and non-random assembly pathways [15]. Functional percolation offers a dynamical mechanism through which such biased exploration and partial persistence can emerge from generic interaction networks, without requiring explicit replication or selection. Whether analogous cascade-driven regimes existed in specific prebiotic chemistries remains an open empirical question.
As shown in Secs. 2–4, functional percolation can be interpreted as a transition in physical interaction networks. When effective connectivity, response thresholds, and coherence times jointly exceed minimal requirements, cascades support persistent, discriminative response patterns rather than isolated or indiscriminate activity.
@@
The present framework does not replace these constraints. Instead, it identifies a complementary, network-level mechanism by which persistence and bias can emerge before replication, heredity, or high-fidelity information storage.
Functional percolation introduces a mesoscopic persistence window: Structured functional responses can recur across realizations through cascade dynamics alone. This provides a dynamical route to biased accumulation in function space, even when microphysical persistence remains limited.
@@ 7 Conclusion
In this work, we have identified a minimal, substrate-agnostic mechanism by which such feasibility can emerge. Thresholded cascade dynamics in connected interaction networks give rise to a connectivity-driven transition—functional percolation—in the accessibility and persistence of structured input–output response mappings. Near this transition, cascades are system-spanning yet discriminative, enabling biased, partially persistent functional organization without requiring long-term memory, replication, or hereditary storage.
The mechanism relies only on generic physical ingredients: interacting units, nonlinear response thresholds, transmission of influence, and finite coherence times. No biochemical specialization, evolutionary optimization, or teleological structure is assumed. Persistence arises over the duration of cascades themselves, allowing functional responses to recur across realizations and introducing directionality and bias into exploration of functional and assembly space.
@@
Acknowledgments and bibliography
  • Acknowledgment of discussions with Henrik Jeldtoft Jensen, Imperial College London

### v2

@@ Abstract
The origin of life poses a problem of combinatorial feasibility: How can temporally supported functional organization arise in exponentially branching assembly spaces when unguided exploration behaves as a memoryless random walk? We show that nonlinear threshold-cascade dynamics in connected interaction networks provide a minimal, substrate-agnostic mechanism that can soften this obstruction. Below a critical connectivity threshold, cascades die out locally and structured input–output response mappings remain sparse and transient—a “functional desert” in which accumulation is dynamically unsupported. Near the critical percolation threshold, system-spanning cascades emerge, enabling discriminative functional responses that recur over finite coherence timescales. We illustrate this transition using a minimal toy model and generalize the argument to arbitrary networked systems. Also near criticality, cascades introduce finite-timescale structural and functional coherence, directional bias, and weak dynamical path-dependence into otherwise memoryless exploration, allowing biased accumulation over short coherence timescales. This connectivity-driven transition—functional percolation—requires only generic ingredients: interacting units, nonlinear thresholds, influence transmission, and non-zero coherence times. The mechanism does not explain specific biochemical pathways, but it identifies a necessary dynamical regime in which structured functional organization can emerge and be temporarily supported over finite coherence intervals, providing a physical foundation for how combinatorial feasibility barriers can be crossed through network dynamics alone.
1 Introduction
@@
Figure 1: Even a microscopic collection of interacting units can support functional accumulation. Minimal interaction structure already supports non-trivial functional responses. Two input nodes AA and $B$ (e.g., coarse-grained stimuli or activation sites) influence a downstream node $C$ through directed couplings. Node $C$ is a thresholded unit with threshold parameter $\phi$: It activates if the weighted fraction of active inputs exceeds $\phi$. For $\phi\leq 0.5$, activation of either input suffices and $C$ implements an OR-like response. For $\phi>0.5$, both inputs are required and $C$ implements an AND-like response. Because the interaction links and response rule remain coherent over the cascade timescale, structured multi-input responses can recur over short coherence intervals within the same dynamical regime, supporting a minimal form of functional accumulation.
A central challenge in the origin of life is the emergence of functionally coherent organization over non-zero timescales in vast combinatorial spaces. Previous work has shown that unguided exploration behaves effectively as a memoryless random walk, making reliable accumulation dynamically unsupported rather than merely improbable [4, 5]. In such spaces, the number of possible molecular or functional configurations grows super-exponentially with system size, rendering reliable accumulation of organized structure vanishingly unlikely without bias or mechanisms that preserve coherence in exploration over finite timescales.
Although motivated by origin-of-life feasibility, the mechanism is stated for generic interacting units (molecules, proteins, genes, cells, neurons, agents), and the conclusions concern a substrate-agnostic transition in the space of realizable input–output response mappings.
Here, coherence refers to the finite duration over which (i) the effective interaction structure remains sufficient for information transmission and (ii) the local response rules remain well-defined, so that a cascade can propagate and implement a reproducible input–output mapping over that timescale. Accumulation refers not to the growth of specific molecular objects, but to a more generalized notion: The emergence and finite-timescale support of structured functional organization, as revealed through non-trivial system responses to environmental inputs. In this sense, our usage aligns with Schrödinger’s qualitative characterization of life as maintaining and amplifying order in far-from-equilibrium conditions. Schrödinger’s emphasis on order referred not to formal entropy reduction, but to the finite-timescale maintenance of organized, functional structure in far-from-equilibrium systems [13].
@@
Thus, while prior work has emphasized the infeasibility of assembling complex structures in vast configuration spaces, the present framework makes clear that the deeper obstruction lies in function space: Below a critical interaction regime, structured input–output mappings cannot remain coherent over non-zero timescales. Therefore the above ’feasibility desert’ becomes a ’functional desert’, in which meaningful functional organization is dynamically unsupported.
This motivates the central question of the present work: Can cascade dynamics provide a minimal physical mechanism that relieves the combinatorial feasibility obstruction? Specifically, we ask whether connectivity-driven transitions in cascade behavior can enable the emergence of structured functional responses that recur over finite coherence timescales in otherwise unguided systems, thereby allowing accumulation in exponentially branching spaces.
@@ 3.1 Toy Example: Minimal Functional Cascades
Imagine three system elements—molecules, neurons, cells, or other interacting units—that come together randomly and establish effective channels of influence among them. In physical systems, such interactions are typically mediated by contact or proximity, so small multi-unit encounters need not be implausible across many domains. If these interactions are sufficient to form a minimally connected component, then activity at one element can, in principle, reach the others. In this informal sense, the system has crossed a basic connectivity threshold analogous to percolation: Local interactions are no longer isolated, but linked into a coherent interaction structure.
@@
Importantly, such functional responses require only finite-timescale coherence. The interaction links must remain intact long enough for influence to propagate, and the response rule at C must remain coherent long enough for inputs to be integrated. No long-term memory, storage, or repeatability is assumed. The coherence required is strictly local in time and limited to the duration of the cascade itself.
In this sense, even a tiny connected system can begin to "cross the functional desert." Rather than requiring the accumulation of specific molecular objects, the system accumulates functional organization: Structured input–output relationships that can, in principle, be reused and extended within recurring coherence intervals. The toy model provides a concrete, intuitive demonstration of how minimal interactions can already support nontrivial functional structure.
3.2 General Mechanism: Cascades and Finite-timescale coherence
The toy example illustrates a more general mechanism that applies to networked systems of arbitrary size. Consider a system composed of many interacting nodes. As connectivity increases, the system undergoes a percolation transition: A giant connected component emerges, allowing influences to propagate across large portions of the network [17]. Random graph models provide a conservative baseline for this transition, demonstrating that cascade-capable connectivity does not require specialized architecture or biological optimization.
@@ 3.2 General Mechanism: Cascades and Finite-timescale coherence
Under these minimal conditions, cascades can integrate multiple inputs into organized system-level responses. Because physical interactions are proximity-mediated, small interaction motifs and short-lived channels of influence need not be implausible in many domains; however, the central requirement here is simply that connectivity and response rules remain coherent over the cascade timescale. In sufficiently connected systems, this is enough for structured input–output response mappings to be realized and to recur over short coherence intervals, without requiring long-term memory, stored programs, or a single combinatorial leap to highly organized assemblies. Functional organization thus becomes accessible directly from connectivity and nonlinear dynamics whenever influence can propagate and thresholds implement multi-input integration. This connectivity-driven transition in the accessibility of functional responses, coinciding with cascade emergence near percolation criticality, is what we refer to as functional percolation, and has been studied elsewhere [17, 18, 19].
@@ 4 Consequences of the General Mechanism
4.1 Finite-timescale coherence
The core argument implies two minimal and distinct forms of coherence. Structural coherence refers to the coherence of interaction connectivity and susceptibility profiles over the timescale of a cascade. For influence to propagate, the network structure must remain intact long enough for signals to traverse multiple nodes [17, 6].
Functional coherence refers to the coherence of response mappings over the realization timescale. For a system to exhibit a structured functional response, the rules governing how nodes integrate inputs must remain stable long enough for inputs to be transformed into organized outputs [18, 14].
Crucially, neither form of coherence assumes long-term stability, stored memory, inheritance, or replication. Coherence here means only that relevant structures remain coherent over the duration of a single realization. This minimal notion of coherence is sufficient to support cascades and functional responses without invoking biological machinery or evolutionary processes [19].
@@ 4.2 Accumulation
In purely memoryless exploration of combinatorial spaces, partial progress is lost between realizations, preventing reliable buildup of organization, much like the "monkeys typing Shakespeare" intuition [3]. Finite-timescale structural and functional coherence prevents this total erasure by allowing coherent response patterns to recur across perturbations.
@@ 4.3 Directionality and Bias
In this framework, directionality refers not to the direction of signal propagation per se, but to the directionality of accumulation in functional and assembly space [5]. Cascades, together with minimal structural and functional coherence, bias which functional organizations are realized and can recur over short coherence timescales.
Because node responses depend on collective inputs, certain response patterns recur more readily than others. This introduces a bias in how structured functional organizations are explored and supported for recurrence over finite coherence times, even in the absence of design, optimization, or selection [17, 12].
@@
Importantly, this directionality is non-teleological. It does not reflect purpose, intention, or evolutionary selection, but arises as a structural consequence of cascades operating in finite-timescale coherent interaction networks [8].
@@ 5 Feasibility and Crossing the Functional Desert
Figure 2: Functional percolation as a feasibility transition in function space (schematic). As effective interaction connectivity increases, threshold-cascade dynamics organize into three regimes. Below the transition (subcritical), cascades die out locally and realizable input–output response functions are sparse and nonpersistent. Near the functional percolation threshold, non-saturating system-spanning cascades emerge, sharply expanding the accessible repertoire of response functions that recur over finite coherence timescales and enabling long-range directed influence. Above the transition (supercritical), increasing connectivity strengthens collective coupling, reducing functional diversity and constraining responses toward more correlated global patterns without eliminating information flow. The near-critical regime constitutes a narrow feasible window in which accumulation in an exponentially branching space becomes possible through finite-timescale coherence and bias, without collapse or saturation of functional distinctions.
@@
Near the critical transition, however, cascades are system-spanning without being fully saturating. This regime supports maximal functional discrimination, long-range directed influence, and the emergence of a rich repertoire of response mappings that recur over finite coherence intervals [Fig. 2] [19, 18]. Only in this narrow window does structured functional organization become both realizable and capable of recurrence across realizations. An ’oasis’ in the functional desert.
Thus, near criticality, cascades bias which functional organizations are realized. Exploration becomes directionally constrained in functional and assembly space, introducing path-dependence in what is built [5, 14]. Accumulation is no longer symmetric or transient, but structured and supported for recurrence over finite coherence timescales. Quantitative characterization of this weak dynamical path-dependence in specific models is left for future work.
In this sense, functional percolation constitutes a feasibility transition. Random exploration in combinatorial spaces is converted into directionally biased accumulation through the emergence of discriminative cascades that remain coherent over finite timescales [5, 14]. Accumulation becomes physically supported, not merely statistically possible.
@@ 6 Discussion
Across these domains, structured functional responses to inputs correspond to collective behaviors with non-zero coherence over cascade/interaction timescales, enabled by interaction topology, response thresholds, and timescale coherence rather than by detailed biochemical specialization. Random graph models provide a conservative baseline for this transition, demonstrating that cascade-enabled functional responses do not rely on specialized structure or biological optimization.
More generally, the bridge from microscopic interaction to macroscopic organization is provided by compositional structure. If small interaction motifs can implement structured responses over their coherence timescales, then those motifs can function as modules that participate in larger-scale networks of influence. As interaction scope increases—through additional contacts, proximity-mediated encounters, or higher effective connectivity—modules can combine to form richer functional organizations without requiring that large assemblies appear all at once. This perspective applies beyond biochemistry: neuronal circuits, cellular signaling modules, ecological subsystems, and social groups all exhibit local interaction motifs that can combine into higher-order organizations. Accordingly, the present framework connects minimal cascade-enabled function to the scalable emergence of organized systems by emphasizing modular composition rather than abrupt large-scale assembly.
Empirical assembly-theory studies emphasize that real chemical systems occupy a highly structured and historically constrained region of object space, characterized by reuse of substructures and non-random assembly pathways [15]. Functional percolation offers a dynamical mechanism through which such biased exploration and partial finite-timescale coherence can emerge from generic interaction networks, without requiring explicit replication or selection. Whether analogous cascade-driven regimes existed in specific prebiotic chemistries remains an open empirical question.
As shown in Secs. 2–4, functional percolation can be interpreted as a transition in physical interaction networks. When effective connectivity, response thresholds, and coherence times jointly exceed minimal requirements, cascades support discriminative response patterns that remain coherent over finite coherence timescales rather than isolated or indiscriminate activity.
@@
The present framework does not replace these constraints. Instead, it identifies a complementary, network-level mechanism by which coherence and bias can emerge before replication, heredity, or high-fidelity information storage.
Functional percolation introduces a mesoscopic coherence window: Structured functional responses can recur across realizations through cascade dynamics alone. This provides a dynamical route to biased accumulation in function space, even when microphysical persistence remains limited.
@@ 7 Conclusion
In this work, we have identified a minimal, substrate-agnostic mechanism by which such feasibility can emerge. Thresholded cascade dynamics in connected interaction networks give rise to a connectivity-driven transition—functional percolation—in the accessibility and finite-timescale coherence of structured input–output response mappings. Near this transition, cascades are system-spanning yet discriminative, enabling biased, partially persistent functional organization without requiring long-term memory, replication, or hereditary storage.
The mechanism relies only on generic physical ingredients: interacting units, nonlinear response thresholds, transmission of influence, and finite coherence times. No biochemical specialization, evolutionary optimization, or teleological structure is assumed. Coherence arises over the duration of cascades themselves, allowing functional responses to recur across realizations and introducing directionality and bias into exploration of functional and assembly space.
@@
Acknowledgments
Acknowledgment of discussions with Henrik Jeldtoft Jensen, Imperial College London.